# TimeDistill: Efficient Long-Term Time Series Forecasting with MLP via Cross-Architecture Distillation

### v1

@@ 4.1 Multi-Scale Distillation
$\mathbf{\hat{Y}}_{x}^{m}=\mathrm{Conv}(\mathbf{\hat{Y}}_{x}^{m-1},\mathrm{stride}=2),$ (3)
@@
$\mathcal{L}_{scale}^{\mathbf{Y}}=\textstyle\sum_{m=0}^{M}||\mathbf{\hat{Y}}_{t}^{m}-\mathbf{\hat{Y}}_{s}^{m}||^{2}/(M+1).$ (4)
@@
$\mathbf{H}^{\prime}_{t}=\text{Regressor}(\mathbf{H}_{t}),$ (5)
@@
$\mathcal{L}_{scale}^{\mathbf{H}}=\textstyle\sum_{m=0}^{M}||\mathbf{H}_{t}^{m}-\mathbf{H}_{s}^{m}||^{2}/(M+1).$ (6)
4.2 Multi-Period Distillation
@@
$\mathbf{A}_{x}=\text{Amp}(\text{FFT}(\mathbf{\hat{Y}}_{x})),$ (7)
@@
To reduce the influence of minor frequencies and avoid noise introduced by less meaningful frequencies (Wu et al., 2022; Zhou et al., 2022), we propose a distribution-based matching scheme. We use a softmax function with a colder temperature to highlight the most significant frequencies:
$\mathbf{Q}_{x}^{\mathbf{Y}}={\exp\bigl{(}\mathbf{A}_{x}^{i}/\tau\bigr{)}}/{\sum\nolimits_{j=1}^{S/2}\exp\bigl{(}\mathbf{A}_{x}^{j}/\tau\bigr{)}},$ (8)
where $\mathbf{Q}_{x}^{\mathbf{Y}}\in\mathbb{R}^{\frac{S}{2}\times C}$ and $\tau$ is a temperature parameter that controls the sharpness of the distribution. We set $\tau=0.5$ by default. The period distribution $\mathbf{Q}_{x}^{\mathbf{Y}}$ represents the multi-period pattern in the prediction time series, which we want the MLP to learn from the teacher. We use KL divergence to match these distributions (Hinton, 2015). We define the multi-period distillation loss at the prediction level as:
$\mathcal{L}_{period}^{\mathbf{Y}}=\text{KL}\bigl{(}\mathbf{Q}_{t}^{\mathbf{Y}},\mathbf{Q}_{s}^{\mathbf{Y}}\bigr{)}.$ (9)
@@
$\mathcal{L}_{period}^{\mathbf{H}}=\text{KL}\bigl{(}\mathbf{Q}_{t}^{\mathbf{H}},\mathbf{Q}_{s}^{\mathbf{H}}\bigr{)}.$ (10)
4.3 Overall Optimization and Theoretical Analaysis
@@
$\mathcal{L}=\mathcal{L}_{sup}+\alpha\cdot\bigl{(}\mathcal{L}_{scale}^{\mathbf{Y}}+\mathcal{L}_{period}^{\mathbf{Y}}\bigr{)}+\beta\cdot\bigl{(}\mathcal{L}_{scale}^{\mathbf{H}}+\mathcal{L}_{period}^{\mathbf{H}}\bigr{)},$ (12)
@@
Theoretical Interpretations. We provide a theoretical understanding of multi-scale and multi-period distillation loss from a novel data augmentation perspective. We further show that the proposed distillation loss can be interpreted as training with augmented samples derived from a special mixup (Zhang, 2017) strategy. The distillation process augments data by blending ground truth with teacher predictions, analogous to label smoothing in classification, and provides several benefits for time series forecasting: (1) Enhanced Generalization: It enhances generalization by exposing the student model to richer supervision signals from augmented samples, thus mitigating overfitting, especially with limited or noisy data. (2) Explicit Integration of Patterns: The augmented supervision signals explicitly incorporate patterns across multiple scales and periods, offering insights that are not immediately evident in the raw ground truth. (3) Stabilized Training Dynamics: The blending of targets softens the supervision signals, which diminishes the model’s sensitivity to noise and leads to more stable training phases. This will in turn support smoother optimization dynamics and fosters improved convergence. For clarity, our discussion is centered at the prediction level. We present the following theorem:
@@
Let $(x,y)$ denote original input data pairs and $(x,y^{t})$ represent corresponding teacher data pairs. Consider a data augmentation function $\mathcal{A}(\cdot)$ applied to $(x,y)$, generating augmented samples $(x^{\prime},y^{\prime})$. Define the training loss on these augmented samples as $\mathcal{L}_{aug}=\sum_{(x^{\prime},y^{\prime})\in\mathcal{A}(x,y)}|f_{s}(x^{\prime})-y^{\prime}|^{2}$. Then, the following inequality holds: $\mathcal{L}_{sup}+\eta\mathcal{L}_{scale}\geq\mathcal{L}_{aug},$ when $\mathcal{A}(\cdot)$ is instantiated as a mixup function (Zhang, 2017) that interpolates between the original input data $(x,y)$ and teacher data $(x,y^{t})$ with a mixing coefficient $\lambda=\frac{1}{1+\eta}$, i.e. $y^{\prime}=\lambda y+(1-\lambda)y^{t}$.
We provide proof of Theorem 4.1 in Appendix B. Theorem 4.1 suggests that optimizing multi-scale distillation loss $\mathcal{L}_{\text{scale}}$ jointly with supervised loss $\mathcal{L}_{\text{sup}}$ is equivalent to minimizing an upper bound on a special mixup augmentation loss. In particular, we mix multi-scale teacher predictions $\{\mathbf{\hat{Y}}_{t}^{(m)}\}_{m=0}^{M}$ with ground truth $\mathbf{Y}$, thereby allowing MLP to learn more informative time series temporal pattern. Similarly, we present a theorem for understanding $\mathcal{L}_{period}$.
@@
The proof can be found in Appendix B. Theorem 4.2 shows that optimizing the multi-period distillation loss $\mathcal{L}_{\text{period}}$ jointly with the supervised loss $\mathcal{L}_{\text{sup}}$ is equivalent to minimizing an upper bound on the KL divergence between the student period distribution $\mathcal{X}(f_{s}(x^{\prime}))$ (or $\mathbf{Q}_{s}$) and a mixed period distribution $y^{\prime}$ (or $\mathbf{Q}_{y}+\lambda\,\mathbf{Q}_{t}$).
5 Experiment
Table 1: Long-term time series forecasting results with prediction lengths $S~{}\in\{96,192,336,720\}$. A lower MSE or MAE indicates a better prediction. For consistency, we maintain a fixed input length of 720 throughout all the experiments. Results are averaged from all prediction lengths. The best performance is highlighted in red, and the second-best is underlined. Full results are listed in Appendix M.
@@ 5 Experiment
(Ours) (2024) (2024) (2024a) (2023) (2023) (2022) (2022) (2021)
@@
5.1 Experimental Setup
Datasets and Baselines. We run experiments to evaluate the performance and efficiency of TimeDistill on 8 widely used benchmarks: Electricity (ECL), the ETT datasets (ETTh1, ETTh2, ETTm1, ETTm2), Solar, Traffic, and Weather, following (Liu et al., 2024; Wang et al., 2024a; Luo & Wang, 2024). To examine the effectiveness of our method across diverse tasks, we compare TimeDistill with 8 baseline models that cover a range of architectures. Specifically, we use Transformer-based models: iTransformer (Liu et al., 2024), PatchTST (Nie et al., 2023), FEDformer (Zhou et al., 2022), Autoformer (Wu et al., 2021); CNN-based models: ModernTCN (Luo & Wang, 2024), MICN (Wang et al., 2023), TimesNet (Wu et al., 2022); and an MLPs-based model: TimeMixer (Wang et al., 2024a). More details can be found in Appendix C.
5.2 Main Results
@@
Efficiency Analysis. Beyond its strong predictive performance, another notable advantage of TimeDistill is its extremely lightweight architecture, as it is simply an MLP. Figure 2 in the Introduction section shows the trade-off between inference time, memory footprint, and performance. We can observe that TimeDistill can achieve up to speedup and up to 130× fewer parameters compared with baselines. This property makes TimeDistill suitable for deployment on devices with limited computational resources and in latency-sensitive applications that require fast inference. Compared to previous Transformer-based method Autoformer, we achieve 196× speedup as shown in Table 10. We list full results of efficiency analysis in Appendix I.
5.3 Versatility of TimeDistill
Table 2: Performance improvement by TimeDistill with different teachers. $\Delta_{MLP}$, $\Delta_{Teacher}$ indicate the improvement of MLP+TimeDistill over a trained MLP and Teacher, respectively. We report average MSE across all prediction lengths. Full results are in Appendix E.
@@ 5.3 Versatility of TimeDistill
Figure 7: Forecasting performance is evaluated with different look-back lengths of $T\in\{96,192,336,720\}$, and the results are averaged across all prediction lengths $S\in\{96,192,336,720\}$ on eight datasets. The complete results are provided in Appendix N.
@@
Different Teachers. We use ModernTCN as the teacher model for our main results. In Table 2, we present results using other teacher models, such as iTransformer, TimeMixer, and PatchTST. Results with additional teachers are provided in Appendix E. We observe that TimeDistill can effectively learn from various teachers, improving MLP by up to 13.87%. Furthermore, TimeDistill achieves significant performance improvements over the teachers themselves, with gains of up to 21.41%. We hypothesize two key reasons for these improvements. First, the student MLP model already demonstrates strong learning capabilities; for instance, on the Solar dataset, MLP outperforms most teachers. Second, the multi-scale and multi-period KD approach delivers both temporal and frequency knowledge from teachers to MLP, offering additional valuable insights. Similar findings have been reported in recent KD studies (Allen-Zhu & Li, 2020; Zhang et al., 2021; Guo et al., 2023), which suggest that integrating diverse views from multiple models can enhance performance.
@@
Different Students. To evaluate whether TimeDistill can enhance the performance of other lightweight models, we select two MLPs-based models, TSMixer (Chen et al., 2023) and LightTS (Zhang et al., 2022), as students. Unlike a simple MLP, these MLPs-based models consist of multiple MLPs and operate in a channel-dependent manner. Consequently, they are generally more complex than a standard MLP, which results in reduced efficiency. We use ModernTCN as the teacher model, consistent with other experiments. As shown in Table 3, TimeDistill consistently improves the performance of TSMixer and LightTS, achieving remarkable MSE reductions of 6.26% and 8.02%, respectively. These results demonstrate that TimeDistill is highly adaptable and can effectively enhance other lightweight methods. We also explore the influence of student MLP architecture (e.g. layer number and hidden dimension) on TimeDistill in Appendix F.
@@
Table 3: Performance improvement of TimeDistill with different students on ETTh1, averaged across all prediction lengths. $\Delta_{Student}$ represents the improvement of Student+TimeDistill over the original student. Additional results are in Appendix F.
Student Models MLP LightTS TSMixer
(2022) (2023)
MSE MAE MSE MAE MSE MAE
Student 0.502 0.489 0.465 0.471 0.471 0.474
+TimeDistill 0.428 0.445 0.436 0.445 0.433 0.446
$\Delta_{Student}$ 14.74% 9.00% 6.26% 5.57% 8.02% 5.92%
@@
Different Look-Back Window Lengths. The length of the look-back window significantly influences forecasting accuracy, as it determines how much historical information can be utilized for learning. Figure 7 presents the average MSE results across all eight datasets. Overall, the performance of all models, particularly MLP, improves as the look-back window size increases. Notably, TimeDistill consistently enhances the performance of MLP and outperforms the teacher models across all look-back window lengths.
5.4 Deeper Analysis into Our Distillation Framework
@@
Table 4: Ablation study measured by MSE on different components of TimeDistill (Teacher: ModernTCN). The best performance is in red, and the second-best is underlined. More ablation study results are listed in Appendix O.
@@
Figure 8: Prediction comparison across temporal scales and spectrograms before and after distillation on ETTh1. MSE for MLP, Teacher (ModernTCN), and TimeDistill are 0.790, 0.365, and 0.366, showing TimeDistill bridges temporal and frequency domain gaps via multi-scale and multi-period distillation. More cases in Appendix K.
Table 5: Win ratio (%) of MLP and TimeDistill vs. ModernTCN under input-720-predict-96 setting. $U_{M}$ and $U_{T}$ denote samples where MLP and TimeDistill outperform the teacher. Win Keep, $\frac{|U_{M}\cap U_{T}|}{|U_{M}|}$, shows TimeDistill’s retention of MLP’s wins.
@@
6 Conclusion and Future work

### v2

@@ 4.1. Multi-Scale Distillation
(3) $\mathbf{\hat{Y}}_{x}^{m}=\mathrm{Conv}(\mathbf{\hat{Y}}_{x}^{m-1},\mathrm{stride}=2),$
@@
[…]
@@
(5) $\mathbf{H}^{\prime}_{t}=\text{Regressor}(\mathbf{H}_{t}),$
@@
[…]
4.2. Multi-Period Distillation
@@
(7) $\mathbf{A}_{x}=\text{Amp}(\text{FFT}(\mathbf{\hat{Y}}_{x})),$
@@
To reduce the influence of minor frequencies and avoid noise introduced by less meaningful frequencies (timesnet; fedformer), we propose a distribution-based matching scheme. We use softmax function with a colder temperature to highlight the most significant frequencies:
(8) $\mathbf{Q}_{x}^{\mathbf{Y}}={\exp\bigl(\mathbf{A}_{x}^{i}/\tau\bigr)}/{\sum\nolimits_{j=1}^{S/2}\exp\bigl(\mathbf{A}_{x}^{j}/\tau\bigr)},$
where $\mathbf{Q}_{x}^{\mathbf{Y}}\in\mathbb{R}^{\frac{S}{2}\times C}$ and $\tau$ is a temperature parameter that controls the sharpness of the distribution. We set $\tau=0.5$ by default. The period distribution $\mathbf{Q}_{x}^{\mathbf{Y}}$ represents the multi-period pattern in the prediction time series, which we want the MLP to learn from the teacher. We use KL divergence to match these distributions (hinton2015distilling). We define multi-period distillation loss at the prediction level as:
(9) $\mathcal{L}_{period}^{\mathbf{Y}}=\text{KL}\bigl(\mathbf{Q}_{t}^{\mathbf{Y}},\mathbf{Q}_{s}^{\mathbf{Y}}\bigr).$
@@
(10) $\mathcal{L}_{period}^{\mathbf{H}}=\text{KL}\bigl(\mathbf{Q}_{t}^{\mathbf{H}},\mathbf{Q}_{s}^{\mathbf{H}}\bigr).$
4.3. Overall Optimization and Theoretical Analysis
@@
(12) $\mathcal{L}=\mathcal{L}_{sup}+\alpha\cdot\bigl(\mathcal{L}_{scale}^{\mathbf{Y}}+\mathcal{L}_{period}^{\mathbf{Y}}\bigr)+\beta\cdot\bigl(\mathcal{L}_{scale}^{\mathbf{H}}+\mathcal{L}_{period}^{\mathbf{H}}\bigr),$
@@
Theoretical Interpretations. We provide a theoretical understanding of multi-scale and multi-period distillation loss from a novel data augmentation perspective. We further show that the proposed distillation loss can be interpreted as training with augmented samples derived from a special mixup (mixup) strategy. The distillation process augments data by blending ground truth with teacher predictions, analogous to label smoothing in classification, and provides several benefits for time series forecasting: (1) Enhanced Generalization: It enhances generalization by exposing the student model to richer supervision signals from augmented samples, thus mitigating overfitting, especially with limited or noisy data. (2) Explicit Integration of Patterns: The augmented supervision signals explicitly incorporate patterns across multiple scales and periods, offering insights that are not immediately evident in the raw ground truth. (3) Stabilized Training Dynamics: The blending of targets softens the supervision signals, which diminishes the model’s sensitivity to noise and leads to more stable training phases. This will in turn support smoother optimization dynamics and fosters improved convergence. For clarity, our discussion is centered at the prediction level. We present the following theorem:
@@
Let $(x,y)$ denote original input data pairs and $(x,y^{t})$ represent corresponding teacher data pairs. Consider a data augmentation function $\mathcal{A}(\cdot)$ applied to $(x,y)$, generating augmented samples $(x^{\prime},y^{\prime})$. Define the training loss on these augmented samples as $\mathcal{L}_{aug}=\sum_{(x^{\prime},y^{\prime})\in\mathcal{A}(x,y)}|f_{s}(x^{\prime})-y^{\prime}|^{2}$. Then, the following inequality holds: $\mathcal{L}_{sup}+\eta\mathcal{L}_{scale}\geq\mathcal{L}_{aug},$ when $\mathcal{A}(\cdot)$ is instantiated as a mixup function (mixup) that interpolates between the original input data $(x,y)$ and teacher data $(x,y^{t})$ with a mixing coefficient $\lambda=\frac{1}{1+\eta}$, i.e. $y^{\prime}=\lambda y+(1-\lambda)y^{t}$.
We provide proof of Theorem 4.1 in Appendix A. Theorem 4.1 suggests that optimizing multi-scale distillation loss $\mathcal{L}_{\text{scale}}$ jointly with supervised loss $\mathcal{L}_{\text{sup}}$ is equivalent to minimizing an upper bound on a special mixup augmentation loss. In particular, we mix multi-scale teacher predictions $\{\mathbf{\hat{Y}}_{t}^{(m)}\}_{m=0}^{M}$ with ground truth $\mathbf{Y}$, allowing MLP to learn more informative time series temporal pattern. Similarly, we present a theorem for understanding $\mathcal{L}_{period}$.
@@
The proof can be found in Appendix A. Theorem 4.2 shows that optimizing the multi-period distillation loss $\mathcal{L}_{\text{period}}$ jointly with the supervised loss $\mathcal{L}_{\text{sup}}$ is equivalent to minimizing an upper bound on the KL divergence between the student period distribution $\mathcal{X}(f_{s}(x^{\prime}))$ (or $\mathbf{Q}_{s}$) and a mixed period distribution $y^{\prime}$ (or $\mathbf{Q}_{y}+\lambda\,\mathbf{Q}_{t}$).
5. Experiment
Table 1. Long-term time series forecasting results with prediction lengths $S~\in\{96,192,336,720\}$. A lower MSE or MAE indicates a better prediction. For consistency, we maintain a fixed input length of 720 throughout all the experiments. Results are averaged from all prediction lengths. The best performance is highlighted in red, and the second-best is underlined. Full results are listed in Appendix R.
@@ 5. Experiment
(Ours) ((itransformer)) ((moderntcn)) ((timemixer)) ((patchtst)) ((micn)) ((fedformer)) ((timesnet)) ((autoformer))
@@
5.1. Experimental Setup
Datasets and Baselines. We run experiments to evaluate the performance and efficiency of TimeDistill on 8 widely used benchmarks: Electricity (ECL), the ETT datasets (ETTh1, ETTh2, ETTm1, ETTm2), Solar, Traffic, and Weather, following (itransformer; timemixer; moderntcn). To examine the effectiveness of our method across diverse tasks, we compare TimeDistill with 8 baseline models that cover a range of architectures. Specifically, we use Transformer-based models: iTransformer (itransformer), PatchTST (patchtst), FEDformer (fedformer), Autoformer (autoformer); CNN-based models: ModernTCN (moderntcn), MICN (micn), TimesNet (timesnet); and an MLPs-based model: TimeMixer (timemixer). More details can be found in Appendix B.
5.2. Main Results
@@
Efficiency Analysis. Beyond its strong predictive performance, another notable advantage of TimeDistill is its extremely lightweight architecture, as it is simply an MLP. Figure 2 in the Introduction section shows the trade-off between inference time, memory footprint, and performance. We can observe that TimeDistill can achieve up to speedup and up to 130× fewer parameters compared with baselines. This property makes TimeDistill suitable for deployment on devices with limited computational resources and in latency-sensitive applications that require fast inference. Compared to previous Transformer-based method Autoformer, we achieve 196× speedup as shown in Table 11. We list full results of efficiency analysis in Appendix J.
5.3. Versatility of TimeDistill
Table 2. Performance improvement by TimeDistill with different teachers. $\Delta_{MLP}$, $\Delta_{Teacher}$ indicate the improvement of MLP+TimeDistill over a trained MLP and Teacher, respectively. We report average MSE across all prediction lengths. Full results are in Appendix F.
@@ 5.3. Versatility of TimeDistill
Figure 7. Forecasting performance is evaluated with different look-back lengths and the results are averaged across all prediction lengths $S\in\{96,192,336,720\}$ on eight datasets. The complete results are provided in Appendix S.
@@
Different Teachers. We use ModernTCN as the teacher model for our main results. In Table 2, we present results using other teacher models, such as iTransformer, TimeMixer, and PatchTST. Results with additional teachers are provided in Appendix F. We observe that TimeDistill can effectively learn from various teachers, improving MLP by up to 13.87%. Furthermore, TimeDistill achieves significant performance improvements over the teachers themselves, with gains of up to 21.41%. We hypothesize two key reasons for these improvements. First, the student MLP model already demonstrates strong learning capabilities; for instance, on the Solar dataset, MLP outperforms most teachers. Second, the multi-scale and multi-period KD approach delivers both temporal and frequency knowledge from teachers to MLP, offering additional valuable insights. Similar findings have been reported in recent KD studies (allen2020towards; zhang2021graph; guo2023linkless), which suggest that integrating diverse views from multiple models can enhance performance.
@@
Different Students. To evaluate whether TimeDistill can enhance the performance of other lightweight models, we select two MLPs-based models, TSMixer (tsmixer), LightTS (lightts), as students. Unlike a simple MLP, these MLPs-based models consist of multiple MLPs and operate in a channel-dependent manner. Consequently, they are generally more complex than a standard MLP, which results in reduced efficiency. Besides, we select a Linear-based model FITS (fits), which is extremely lightweight. We use ModernTCN as the teacher model, consistent with other experiments. As shown in Table 3, TimeDistill consistently improves the performance of TSMixer, LightTS, and FITS, achieving remarkable MSE reductions of 6.26%, 8.02%, and 3.96%, respectively. These results demonstrate that TimeDistill is highly adaptable and can effectively enhance other lightweight methods. We also explore the influence of student MLP architecture (e.g. layer number and hidden dimension) on TimeDistill in Appendix G.
@@
Table 3. Performance improvement of TimeDistill with different students on ETTh1, averaged across all prediction lengths. $\Delta_{Student}$ represents the improvement of Student+TimeDistill over the original student. Additional results are in Appendix G.
Student Models MLP LightTS TSMixer FITS
MSE MAE MSE MAE MSE MAE MSE MAE
Student 0.502 0.489 0.465 0.471 0.471 0.474 0.428 0.443
+TimeDistill 0.428 0.445 0.436 0.445 0.433 0.446 0.411 0.430
$\Delta_{Student}$ 14.74% 9.00% 6.26% 5.57% 8.02% 5.92% 3.96% 2.75%
@@
Different Look-Back Window Lengths. The length of the look-back window significantly influences forecasting accuracy, as it determines how much historical information can be utilized for learning. Figure 7 presents the average MSE results across all eight datasets. Overall, the performance of models in the figure, particularly MLP, generally improves as the look-back window size increases. Notably, TimeDistill consistently enhances the performance of MLP and outperforms the teacher models across all look-back window lengths.
5.4. Deeper Analysis into Our Distillation Framework
@@
Table 4. Ablation study measured by MSE on different components of TimeDistill (Teacher: ModernTCN). The best performance is in red, and the second-best is underlined. More ablation study results are listed in Appendix T.
@@
Figure 8. Prediction comparison across temporal scales and spectrograms before and after distillation on ETTh1. MSE for MLP, Teacher (ModernTCN), and TimeDistill are 0.790, 0.365, and 0.366, showing TimeDistill bridges temporal and frequency domain gaps via multi-scale and multi-period distillation. More cases in Appendix K.
Table 5. Win ratio (%) of MLP and TimeDistill vs. ModernTCN under input-720-predict-96 setting. $U_{M}$ and $U_{T}$ denote samples where MLP and TimeDistill outperform the teacher. Win Keep, $\frac{|U_{M}\cap U_{T}|}{|U_{M}|}$, shows TimeDistill’s retention of MLP’s wins.
@@
Figure 9. Multi-Variate Person Correlation Visualization on ETTh1 and ETTh2 datasets.
Does TimeDistill Learn Multi-Variate Correlation via Distillation? Some teacher models, such as iTransformer and ModerTCN, are designed in a channel-dependent (CD) manner. While this approach can utilize multi-variate correlation effectively, it incurs a higher computational burden. In contrast, advanced methods like PatchTST, TimeMixer, and our method TimeDistill are employed in a channel-independent (CI) manner, which offers better computational efficiency. This leads to an intriguing question: Can TimeDistill enable MLP to learn multi-variate correlation from the teacher without an explicit distillation design? To explore this, we present a preliminary study using Figure 9, which illustrates two cases from the ETTh1 and ETTh2 datasets, each containing seven variables. Our observations reveal that TimeDistill’s multi-variate correlation more closely resembles that of the teacher compared to the MLP model before distillation. This suggests that TimeDistill facilitates learning underlying structures among multi-variates through distillation even though the student MLP is a CI model. These findings highlight a potential benefit of distillation that merits further exploration. We leave this for further work.
6. Conclusion and Future work